OpenStudy (falconmaster52):

What is the rational number equivalent to 3 point 12 with a bar over 12? 3 and 4 over 33 3 and 8 over 33 3 and 10 over 39 3 and 5 over 39

OpenStudy (kittiwitti1):

This?

OpenStudy (falconmaster52):

i think 1 sec

OpenStudy (falconmaster52):

ya

OpenStudy (falconmaster52):

kk

OpenStudy (rperry309):

a suh dude

zepdrix (zepdrix):

Let x=3.1212... Multiply by 100 (to move one full set of the repeated pattern to the other side of the decimal), 100x = 312.1212... Subtract the first number from the second, 100x - x = 312.1212... - 3.1212... Notice both numbers have the same decimal value, so that will subtract off nicely, 100x - x = 312 - 3 Finish it up from there :D

OpenStudy (falconmaster52):

THX

OpenStudy (falconmaster52):

i mean it so dont worry